OpenStudy (jenniferjuice):

Choose the equation below that represents the line that passes through the point (-2, 6) and has a slope of 4.

OpenStudy (jenniferjuice):

y - 2 = 4(x + 6) y + 6 = 4(x - 2) y + 2 = 4(x - 6) y - 6 = 4(x + 2)

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope (m) = 4 (-2,6) -- x1 = -2 and y1 = 6 now sub you info in

OpenStudy (jenniferjuice):

y-6=4(x- -2)

OpenStudy (jenniferjuice):

y-6=4(x+2)

OpenStudy (jenniferjuice):

like that first right?

OpenStudy (texaschic101):

you got it

OpenStudy (texaschic101):

y - 6 = 4(x + 2) is correct

OpenStudy (jenniferjuice):

oh okay thanks :)